the steamboat geyser in yellowstone national park is capable of shooting its hot water up the ground with a speed of 48.0 m/s. How high can this geyser shoot? 


1
Expert's answer
2021-09-23T17:59:06-0400

Highest height, "H=\frac{u^2}{2g}=\frac{-(48)^2}{2\times 9.8}=-117.55m"